# Newton’s method for nonlinear mappings into vector bundles

### v1

@@ 5.3. Local superlinear convergence
In this section we want to prove the local superlinear convergence of Newton’s method. In the spirit of Deuflhard [9], our proof will not rely on quantities, which may allow a qualitative a-priori convergence result, but rather on quantities for which good a-posteriori algorithmic estimates are accessible. This will be the basis of an affine covariant damping strategy, elaborated below. Nevertheless, we will also show some results which allow an a-priori analysis.
Let $V^{-1}_{y}\in\Gamma(\mathcal{L}(\mathcal{E},E_{y}))$ be a vector back-transport. Let $Q:T\mathcal{E}\to\mathcal{E}$ be a connection map and $R:T\mathcal{X}\to\mathcal{X}$ be a retraction. For a fixed $x\in\mathcal{X}$ we can locally perform pullbacks
$R_{x}^{-1}:\mathcal{X}$ $\to T_{x}\mathcal{M}$
$\hat{x}$ $\mapsto\hat{\textbf{x}}:=R_{x}^{-1}(\hat{x})$
@@
$\theta_{z}({x}):=\frac{\big{\|}(Q_{F(x)}\circ F^{\prime}(x))^{-1}\left[(Q_{F(x)}\circ F^{\prime}(x))(\textbf{x}-\textbf{z})-(F(x)-V_{y(x)}^{-1}(y(z))F(z))\right]\big{\|}_{x}}{\|\textbf{x}-\textbf{z}\|_{x}}.$
@@
Let $\mathcal{X}$ be a manifold and $p:\mathcal{E}\to\mathcal{M}$ a vector bundle. Consider a mapping $F:\mathcal{X}\to\mathcal{E}$ and a section $F^{\prime}\in\Gamma(\mathcal{L}(T\mathcal{X},F^{*}T\mathcal{E}))$. Let $Q:T\mathcal{E}\to\mathcal{E}$ be a connection map and $R:T\mathcal{X}\to\mathcal{X}$ be a retraction. Let $x_{\star}\in\mathcal{X}$ be a zero of $F$ and assume that all Newton steps $x_{k}$ are well defined. Assume that
@@
$\|\textbf{x}_{+}-\textbf{x}_{\star}\|_{x}=\|(Q_{F(x)}\circ F^{\prime}(x))^{-1}[(Q_{F(x)}\circ F^{\prime}(x))(\textbf{x}-\textbf{x}_{\star})-(F(x)-V^{-1}_{y(x)}(y(x_{\star}))F(x_{\star}))]\|_{x}.$
@@
Consider a neighborhood $V$ of $x_{\star}$, where Prop. 5.8 holds for some $01>\varepsilon>$ 0. Since $\lim_{x\to x_{\star}}\theta_{x_{\star}}(x)=0$, we find a metric ball $r\mathcal{B}_{x_{\star}}=\{x\in\mathcal{X}:d_{\mathcal{X}}(x,x_{\star})<r\}\subset V$ of radius $r>0$, such that
@@
We now also want to give conditions for local superlinear convergence which can be used for a-priori analysis. With the help of a local norm on $\mathcal{X}$ and a fibrewise norm on $\mathcal{E}$ we can define a norm of a linear operator $A:E_{y}\to T_{x}\mathcal{X}$:
@@
$\|(Q_{F(x)}\circ F^{\prime}(x))^{-1}\|_{E_{y}\rightarrow T_{x}\mathcal{X}}\leq\beta.$
@@
Let $\varepsilon\in(0,1)$. Using that the operator norm of $(Q_{F(x)}\circ F^{\prime}(x))^{-1}$ is bounded and applying Prop. 5.8, we can estimate $\theta_{{x}_{\star}}({x})$ in a neighborhood $V$ of $x_{\star}$ as follows:
$\theta_{x_{\star}}(x)$ $\leq\beta\cdot(1+\varepsilon)\cdot\frac{\|(Q_{F(x)}\circ F^{\prime}(x))(R_{x}^{-1}(x)-R_{x}^{-1}(x_{\star}))-(F(x)-V_{y}^{-1}(y_{\star})F(x_{\star}))\|_{E_{y}}}{d_{\mathcal{X}}(x,x_{\star})}.$
Since $F(x_{\star})=0$ and $F$ is Newton-differentiable at $x_{\star}$, we can apply Prop. 5.9 (i) and obtain $\underset{x\to x_{\star}}{\lim}\theta_{x_{\star}}(x)=0$.
@@ 5.4. Monitoring local convergence
Let us now define a computable quantity to monitor local convergence. For this, we set $x=x_{k}$, $y:=y(x_{k})=p(F(x_{k})),\;y_{\star}:=y(x_{\star})=p(F(x_{\star}))$ and consider the affine-covariant quantity $\theta_{x_{\star}}(x)$ that we have seen in the proof of the local convergence:
@@
Since the target point $x_{\star}$ is not available, we replace $x_{\star}$ by the next iterate $x_{+}:=x_{k+1}$, correspondingly the pullback $\textbf{x}_{\star}$ by $\textbf{x}_{+}=\delta x\in T_{x}{\mathcal{X}}$, and $y_{\star}$ by $y_{+}=p(F(x_{+}))$. By using $\textbf{x}=0_{x}$ and the definition of the $k$-th Newton direction $\delta x=\textbf{x}_{+}-\textbf{x}$, we obtain:
@@
Note, that a vector back-transport is needed here since $Q_{F(x)}\circ F^{\prime}(x)\overline{\delta x_{+}}$ and $F(x_{+})$ do not lie in the same fibre. Thus, we can rewrite $\theta_{x_{+}}(x)$ as
@@
Using that the operator norm of $(Q_{F(x)}\circ F^{\prime}(x))^{-1}$ is bounded and applying Prop. 5.8, we can estimate $\theta_{{x}_{+}}({x})$ as follows:
@@
$\quad+\|(Q_{F(x)}\circ F^{\prime}(x))(R_{x}^{-1}(x_{\star})-R_{x}^{-1}(x_{+}))-(V_{y}^{-1}(y_{\star})F(x_{\star})-V_{y}^{-1}(y_{+})F(x_{+}))\|_{E_{y}}$
Since $F$ is Newton-differentiable at $x_{\star}$ and $F(x_{\star})=0$, we can apply by Prop. 5.9 (i) and obtain
@@
Since $V_{y}^{-1}$ is differentiable, $F^{\prime}$, $F$ and $V_{y}^{-1}$ are locally bounded and $F$ and $y$ are locally Lipschitz continuous, we can estimate the second summand analogously to the proof of Prop. 5.9 by
@@ 6. An Affine Covariant Damping Strategy
In the following we want to develop a method that can also deal with initial values $x_{0}\in\mathcal{X}$ that are not sufficiently close to the solution. Especially in Newton’s method applied in optimization on Riemannian manifolds some strategies including BFGS, Levenberg-Marquardt and trust region methods were introduced, see e.g. [4, 8, 24]. In a linear setting it is well-known that damped Newton methods are one way to globalize the convergence of Newton’s method ([4, 8, 11, 13, 19]). Here, the step size of the Newton direction is scaled by a factor $\lambda\in(0,1]$. Bortoli et. al. provided a damped Newton’s method to solve the problem of finding singularities of vector fields defined on Riemannian manifolds in [5, 6], where A damping factor $\lambda$ is chosen by a linear line search together with a merit function.
We will also use a damping strategy to globalize our method. The choice of the damping factor is more geometrically motivated here and follows an affine covariant strategy in the spirit of Deuflhard [9]. Since we will not use any Riemannian metric, our way of globalization also works for non-Riemannian manifolds. To prepare the choice of our damping factor, we first define a so called Newton path. Our strategy will then be to follow this path by suitably scaling down the Newton directions until the local convergence area of Newton’s method is reached.
6.1. Newton path
First, we define an algebraic Newton path, a generalization of the Newton path in linear spaces [9, Sec. 3.1.4]. Denote by $x(0)$ the starting point of the path. For $\lambda\in[0,1]$ and $x(\lambda)\in\mathcal{X}$ we denote by $y(\lambda):=p(F(x(\lambda)))$ the base point of $F(x(\lambda))$ on the base manifold $\mathcal{M}$ of $\mathcal{E}$. Consider a vector back-transport $V_{y}^{-1}\in\Gamma(\mathcal{L}(\mathcal{E},E_{y}))$. We consider the Newton path problem, which is based on the idea of scaling down the residual by a factor of $1-\lambda$:
@@ 6.1. Newton path
In contrast to the case of linear spaces, a vector back-transport is needed here since $F(x(\lambda))$ and $F(x(0))$ do not lie in the same fibre. The vector back-transport allows us to formulate the Newton path problem in the fixed fibre $E_{y(0)}$, i.e. on a linear space. We call $x(\lambda)$ the algebraic Newton path starting at $x(0)$.
@@
For a mapping $g:\mathbb{X}\rightarrow\mathbb{Y}$ between linear spaces $\mathbb{X}$ and $\mathbb{Y}$ the Newton path is often defined as the solution of the ordinary differential equation [9, Eq. (3.24)]:
$\frac{dx_{d}(\lambda)}{d\lambda}=-g^{\prime}(x_{d}(\lambda))^{-1}g(x_{d}(0)).$
@@
In contrast to the case of linear spaces the algebraic and the differential Newton path do not coincide in general.
@@
Assume that $F:\mathcal{X}\to\mathcal{E}$ is strictly differentiable at $x(0)$ and
$\|({\langle{V_{y(0)}^{-1}}\rangle}^{\prime}(F(x(0)))F^{\prime}(x(0)))^{-1}\|_{E_{y(0)}\to T_{x(0)}\mathcal{X}}$
is finite. Then for sufficiently small $\lambda$ there exists a solution $x(\lambda)$ of the Newton path problem (21) and the simplified Newton method locally converges to this solution with a linear rate.
@@
Application of the implicit function theorem [26], which also holds for strictly differentiable mappings (cf. e.g. [21, Chap. 25]), on
@@
yields the existence of $x(\lambda)$ for sufficiently small $\lambda$ and convergence of the simplified Newton steps towards $x(\lambda)$. ∎
@@
By using the chain rule, we obtain the following expression for the computation of simplified Newton directions $\overline{\delta x^{\lambda}}\in T_{x}\mathcal{X}$ at an initial guess $x\in\mathcal{X}$:
@@
Using $V_{y(0)}^{-1}(y(0))=Id_{E_{y(0)}}$, we calculate the first (simplified) Newton direction $\delta x^{\lambda}$ for the Newton path problem as a solution of the following equation:
@@
Let $V^{-1}_{y}\in\Gamma(\mathcal{L}(\mathcal{E},E_{y}))$ be consistent with a linear connection map $Q:T\mathcal{E}\to\mathcal{E}$. Then the damped Newton direction $\lambda\delta x$ at $x$ of the original problem conicides with the first Newton direction of the Newton path problem and is tangent to the algebraic and to the differential Newton path starting at $x(0)=x$. In particular, we obtain
@@ 6.2. Computation of Damping Factors
Lemma 6.4 shows that if we choose the vector back-transport and the connection map consistent, the first Newton direction for the Newton path problem coincides with the Newton direction for the original problem scaled by a damping factor $\lambda\in[0,1]$. If we successively choose the current iterate $x=x_{k}$ for the starting point $x(0)$ of the Newton path, we obtain a damped Newton method. More precisely, for a previously computed Newton direction and a damping factor $\lambda$ we compute the Newton step
@@
For the choice of the damping factor we again consider our Newton path problem (21). The idea is to choose $\lambda$ sufficiently small such that the simplified Newton method for solving (21) with starting point $x(0)=x$ is likely to converge to the target point $x(\lambda)$ on the Newton path. To monitor the local convergence, we use the estimator $\theta_{x_{+}}({x})$ at the current iterate $x$ for $\theta_{{x}(\lambda)}({x})$. According to (18) and using (24), we have to compute
$\theta_{x_{+}}({x})=\theta_{R_{x}(\lambda\delta x)}({x})=\frac{\|\overline{\delta x_{+}^{\lambda}}\|_{x}}{\|\lambda\delta x\|_{x}}.$
Thus, we need the next simplified Newton direction $\overline{\delta x_{+}^{\lambda}}$ for the Newton path problem (21) at initial guess $x$. This direction is given by the solution of the equation
${\langle{V_{y(x)}^{-1}}\rangle}^{\prime}(F(x))F^{\prime}(x)\overline{\delta x^{\lambda}_{+}}+V^{-1}_{y(x)}(y(x_{+}))F(x_{+})-(1-\lambda)F(x)=0.$
These calculations are repeated with decreasing damping factor until $\theta_{x_{+}}({x})<\Theta_{acc}$ for some acceptable contraction $\Theta_{acc}$ holds. To achieve this, the damping factor $\lambda$ is readjusted iteratively in the following way. If $\theta_{x_{+}}({x})\geq\Theta_{acc}$, choose $\lambda_{+}\in(0,1]$ such that for user defined parameters $0<\Theta_{des}<\Theta_{acc}$ the inequality $\frac{\lambda_{+}}{\lambda}\cdot\theta_{x_{+}}({x})\leq\Theta_{des}$ holds, i.e. we compute the next damping factor by
@@
We summarise the results in an algorithm, which is a modification of [9, Alg. NLEQ-ERR]:
@@
and thus the inner loop terminates after finitely many iterations.
@@
If $\theta_{R_{x}(\lambda\delta x)}(x)>\Theta_{acc}$, we obtain
@@
Since $\Theta_{des}<\Theta_{acc}$, this yields $\lambda_{+}<C\cdot\lambda\text{ with }C<1$. Now choose $\widehat{\lambda}>0$ sufficiently small such that the Newton path with starting point $x(0)=x$ exists (Prop. 6.2). Since
$\theta_{R_{x}(\lambda\delta x)}({x})=\frac{\|\overline{\delta x_{+}^{\lambda}}\|_{x}}{\|\delta x^{\lambda}\|_{x}}.$
measures the ratio between the first and second simplified Newton step for the Newton path problem and the simplified Newton method converges locally to $x(\lambda)$, we obtain
$\theta_{R_{x}(\lambda\delta x)}(x)\leq\Theta_{acc}\;\forall\lambda\leq\widehat{\lambda}.$
@@
By the choice of $\Theta_{des}\in]0,1[$ it is possible to adjust how aggressive the step size strategy acts. If $\Theta_{des}\approx 0$, then Newton’s method follows the Newton path very closely and takes short steps. If $\Theta_{des}\approx 1$, then larger steps will be taken, but the methods might perform less robustly for highly nonlinear problems. The choices $\Theta_{des}=0.5$ and $\Theta_{acc}=1.1\Theta_{des}$ work well in practice.

### v2

@@ 5.3. Local superlinear convergence
In this section we want to prove the local superlinear convergence of Newton’s method. In the spirit of Deuflhard [9], our proof will rely on quantities, which not only allow a qualitative a-priori convergence result, but also for which good a-posteriori algorithmic estimates are accessible. This second aspect will be the basis of an affine covariant damping strategy, elaborated in the next section. In contrast to [9], which relies on affine covariant Lipschitz constants, we take a more qualitative approach and consider affine covariant versions of Newton- and strict differentiabililty, as established in Proposition 5.10.
Consider again $F:\mathcal{X}\to\mathcal{E}$ where the $C^{1}$-Banach manifold $\mathcal{X}$ is equipped with a continuous local norm $\|\cdot\|_{x}$ and an induced metric $d_{\mathcal{X}}$. Let $V^{-1}_{y}\in\Gamma(\mathcal{L}(\mathcal{E},E_{y}))$ be a vector back-transport, $Q\in\Gamma(\mathcal{L}(T\mathcal{E},p^{*}\mathcal{E}))$ be a connection map, and $R:T\mathcal{X}\to\mathcal{X}$ be a $C^{1}$-retraction. For a fixed $x\in\mathcal{X}$ we can locally perform pullbacks
$R_{x}^{-1}:\mathcal{X}$ $\to T_{x}\mathcal{X}$
$\hat{x}$ $\mapsto\hat{\textbf{x}}:=R_{x}^{-1}(\hat{x}).$
@@
$\theta_{z}({x}):=\frac{\big{\|}(Q_{F(x)}\circ F^{\prime}(x))^{-1}\big{[}(Q_{F(x)}\circ F^{\prime}(x))(\textbf{x}-\textbf{z})-(F(x)-V_{y(x)}^{-1}(y(z))F(z))\big{]}\big{\|}_{x}}{\|\textbf{x}-\textbf{z}\|_{x}}.$
@@
Let $\mathcal{X}$ be a manifold and $p:\mathcal{E}\to\mathcal{Y}$ a vector bundle. Consider a mapping $F:\mathcal{X}\to\mathcal{E}$ and a section $F^{\prime}\in\Gamma(\mathcal{L}(T\mathcal{X},F^{*}T\mathcal{E}))$. Let $Q\in\Gamma(\mathcal{L}(T\mathcal{E},p^{*}\mathcal{E}))$ be a connection map and $R:T\mathcal{X}\to\mathcal{X}$ be a retraction. Let $x_{\star}\in\mathcal{X}$ be a zero of $F$ and assume that all Newton steps $x_{k}$ are well defined. Assume that
@@
$\|\textbf{x}_{+}-\textbf{x}_{\star}\|_{x}=\|(Q_{F(x)}\circ F^{\prime}(x))^{-1}\big{[}(Q_{F(x)}\circ F^{\prime}(x))(\textbf{x}-\textbf{x}_{\star})-(F(x)-V^{-1}_{y(x)}(y(x_{\star}))F(x_{\star}))\big{]}\|_{x}.$
@@
Consider a neighborhood $V$ of $x_{\star}$, where Prop. 5.9 holds for some $\varepsilon>0$. Since $\lim_{x\to x_{\star}}\theta_{x_{\star}}(x)=0$, we find a metric ball $r\mathcal{B}_{x_{\star}}=\{x\in\mathcal{X}:d_{\mathcal{X}}(x,x_{\star})<r\}\subset V$ of radius $r>0$, such that
@@
We now also want to give conditions for local superlinear convergence which can be used for a-priori analysis. With the help of a local norm on $\mathcal{X}$ and a fibrewise norm on $\mathcal{E}$ we can use the standard norm of a linear operator $A:E_{y}\to T_{x}\mathcal{X}$:
@@
$\|(Q_{F(x)}\circ F^{\prime}(x))^{-1}\|_{E_{y(x)}\rightarrow T_{x}\mathcal{X}}\leq\beta.$
@@
Let $\varepsilon\in(0,1)$. Using that the operator norm of $(Q_{F(x)}\circ F^{\prime}(x))^{-1}$ is bounded and applying Prop. 5.9, we can estimate $\theta_{{x}_{\star}}({x})$ in a neighborhood $V$ of $x_{\star}$ as follows:
$\theta_{x_{\star}}(x)$ $\leq\beta\cdot(1+\varepsilon)\cdot\frac{\|(Q_{F(x)}\circ F^{\prime}(x))(R_{x}^{-1}(x)-R_{x}^{-1}(x_{\star}))-(F(x)-V_{y}^{-1}(y_{\star})F(x_{\star}))\|_{E_{y(x)}}}{d_{\mathcal{X}}(x,x_{\star})}.$
Since $F(x_{\star})=0$ and $F$ is Newton-differentiable at $x_{\star}$, we can apply Prop. 5.10 (i) and obtain $\underset{x\to x_{\star}}{\lim}\theta_{x_{\star}}(x)=0$.
Observe that the quantities, used in Proposition 5.12 depend on the choice of a norm on $E_{y}$, while the quantities in Proposition 5.11 do not. This suggests that these norms are not strictly necessary to characterize local convergence, one of the core ideas of affine covariant analysis of Newton methods.
@@ 5.4. Monitoring local convergence
Let us now define a computable quantity to monitor local convergence. For this, we set $x=x_{k}$, $y:=y(x)=p(F(x)),\;y_{\star}:=y(x_{\star})=p(F(x_{\star}))$ and consider the affine-covariant quantity $\theta_{x_{\star}}(x)$ that we have seen in the proof of the local convergence:
@@
Since the target point $x_{\star}$ is computationally not available, we replace $x_{\star}$ by the next iterate $x_{+}:=x_{k+1}$, correspondingly the pullback $\textbf{x}_{\star}$ by $\textbf{x}_{+}=\delta x\in T_{x}{\mathcal{X}}$, and $y_{\star}$ by $y_{+}=p(F(x_{+}))$. By using $\textbf{x}=0_{x}$ and the definition of the Newton direction $\delta x=\textbf{x}_{+}-\textbf{x}=(Q_{F(x)}\circ F^{\prime}(x))^{-1}(-F(x))$, we obtain:
@@
A vector back-transport is needed here since $Q_{F(x)}\circ F^{\prime}(x)\overline{\delta x_{+}}$ and $F(x_{+})$ do not lie in the same fibre. Thus, we can rewrite $\theta_{x_{+}}(x)$ as
@@
The following lemma shows that we can use this to detect local convergence.
@@
Using that the operator norm of $(Q_{F(x)}\circ F^{\prime}(x))^{-1}$ is bounded and applying Prop. 5.9, we can estimate $\theta_{{x}_{+}}({x})$ as follows:
@@
$\quad+\|(Q_{F(x)}\circ F^{\prime}(x))(R_{x}^{-1}(x_{\star})-R_{x}^{-1}(x_{+}))-(V_{y}^{-1}(y_{\star})F(x_{\star})-V_{y}^{-1}(y_{+})F(x_{+}))\|_{E_{y}}.$
Since $F$ is Newton-differentiable at $x_{\star}$ and $F(x_{\star})=0$, we can apply by Prop. 5.10 (i) and obtain
@@
Since $V_{y}^{-1}$ and $Q_{F(x)}\circ F^{\prime}(x)$ are locally bounded and $F$ is locally Lipschitz continuous, we can estimate the second summand by using the triangle inequality, $F(x_{\star})=0$ and Prop. 5.9 as
@@ 6. An Affine Covariant Damping Strategy
To obtain convergence from remote initial guesses, Newton’s method is typically equipped with a globalization strategy. Especially in Newton’s method applied in optimization on Riemannian manifolds some strategies including BFGS, Levenberg-Marquardt and trust region methods were introduced, see e.g. [4, 8, 26]. In a linear setting it is well-known that damped Newton methods are one way to globalize the convergence of Newton’s method ([4, 8, 12, 14, 20]). Here, the step size of the Newton direction is scaled by a factor $\lambda\in(0,1]$. Bortoli et. al. provided a damped Newton’s method to solve the problem of finding zeros of vector fields defined on Riemannian manifolds in [5, 6], where a damping factor $\lambda$ is chosen by a linear line search with respect to a merit function.
We will also use a damping strategy to globalize our method. The choice of the damping factor is more geometrically motivated here and follows an affine covariant strategy in the spirit of Deuflhard [9]. Our way of globalization works for a manifold $\mathcal{X}$, equipped with the Banach-type metrics from Section 5.1. In our algorithm we only have to evaluate local norms $\|\cdot\|_{x}$ on tangent spaces, the evalution of $d_{\mathcal{X}}$ and also of norms $\|\cdot\|_{E_{y}}$ on $\mathcal{E}$ is not needed.
To prepare the choice of our damping factor, we first define a so called Newton path. Our strategy will then be to follow this path by suitably scaling down the Newton directions until the local convergence area of Newton’s method is reached. It will turn out that strict differentiability is an appropriate condition to render this strategy well defined.
6.1. Newton paths on Manifolds
Let us define an algebraic Newton path for $F:\mathcal{X}\to\mathcal{E}$ from a manifold $\mathcal{X}$ to a vector bundle $p:\mathcal{E}\to\mathcal{Y}$, a generalization of the Newton path in linear spaces [9, Sec. 3.1.4]. Denote by $x(0)$ the starting point of the path. For $\lambda\in[0,1]$ and $x(\lambda)\in\mathcal{X}$ we denote by $y(\lambda):=p(F(x(\lambda)))\in\mathcal{Y}$ the base point of $F(x(\lambda))$. Consider a vector back-transport $V_{y}^{-1}\in\Gamma(\mathcal{L}(\mathcal{E},E_{y}))$. We consider the Newton path problem, which is based on the idea of scaling down the residual by a factor of $1-\lambda$:
@@ 6.1. Newton paths on Manifolds
In contrast to the case of linear spaces, we need a vector back-transport on the left-hand side since $F(x(\lambda))$ and $F(x(0))$ do not lie in the same fibre. The vector back-transport allows us to formulate the Newton path problem in the fixed fibre $E_{y(0)}$, i.e. on a linear space. We call the mapping $\lambda\to x(\lambda)$ (where it is defined) the algebraic Newton path starting at $x(0)$.
@@
For a mapping $G:\mathbb{X}\rightarrow\mathbb{Y}$ between linear spaces $\mathbb{X}$ and $\mathbb{Y}$ the Newton path can be defined alternatively as the solution of the ordinary differential equation [9, Eq. (3.24)]:
$G^{\prime}(x_{d}(\lambda))x_{d}^{\prime}(\lambda)+G(x_{d}(0))=0.$
@@
In contrast to the case of linear spaces the algebraic and the differential Newton path do not coincide in our more general setting.
@@
Let $F:\mathcal{X}\to\mathcal{E}$ be strictly differentiable at $x(0)$ and ${\langle{V_{y(0)}^{-1}}\rangle}^{\prime}(F(x(0)))F^{\prime}(x(0))\in L(E_{y(0)},T_{x(0)}\mathcal{X})$ continuously invertible. Then for sufficiently small $\lambda$ there exists a solution $x(\lambda)$ of the Newton path problem (28).
@@
Application of the implicit function theorem [28], which also holds for strictly differentiable mappings (cf. e.g. [23, Chap. 25]), on
@@
yields the existence of $x(\lambda)$ for sufficiently small $\lambda$. ∎
@@
By using the chain rule, we obtain the following expression for the computation of simplified Newton directions $\overline{\delta x^{\lambda}}\in T_{x}\mathcal{X}$ at a current iterate $x\in\mathcal{X}$:
@@
Using $V_{y(0)}^{-1}(y(0))=Id_{E_{y(0)}}$, we calculate the first (simplified) Newton direction $\delta x^{\lambda}$ at $x=x(0)$ for the Newton path problem as a solution of the following equation:
@@
Let $x\in\mathcal{X}$, $Q\in\Gamma(\mathcal{L}(T\mathcal{E},p^{*}\mathcal{E}))$ be a linear connection map, and $\lambda\in(0,1]$. Let $\delta x$ be the current Newton direction, given by solving (6).
Consider the Newton path problem (28) with vector transport $V^{-1}_{y}\in\Gamma(\mathcal{L}(\mathcal{E},E_{y}))$ and assume that $V_{y}^{-1}$ is consistent with $Q$ at $e=F(x)$. Then the damped Newton direction $\lambda\delta x$ coincides with the first Newton direction $\delta x^{\lambda}$ of the Newton path problem (28). Moreover, it is tangent to the algebraic and to the differential Newton path starting at $x(0)=x$. In particular, we obtain
@@ 6.2. Computation of Damping Factors
Our damping strategy is based on following Newton paths. At the current iterate $x$ we consider the algebraic Newton path starting at $x(0)=x$. Lemma 6.4 yields the crucial observation that the damped Newton direction is tangential to the Newton path, as along as we choose the vector back-transport and the connection map consistently. Thus for small damping factors, the deviation of the damped Newton step from the Newton path is very small. In our case the damped Newton step is computed from the Newton direction $\delta x$ and a damping factor $\lambda\in(0,1]$ via a retraction:
@@
The idea is now to choose $\lambda$ sufficiently small such that the simplified Newton method for solving (28) with starting point $x(0)=x$ is likely to converge to the target point $x(\lambda)$ on the Newton path, i.e., solves the Newton path problem. By Lemma 6.4 the damped Newton step coincides with the first Newton step of the Newton path problem. Thus we may use the convergence monitor $\theta_{x_{+}}({x})$ from Section 5.4 at the current iterate $x$ to detect local convergence towards the solution of the Newton path problem. According to (25) and using (31), our convergence monitor reads
$\theta_{x_{+}}({x})=\theta_{R_{x}(\lambda\delta x)}({x})=\frac{\|\overline{\delta x_{+}^{\lambda}}\|_{x}}{\|\lambda\delta x\|_{x}}.$ (32)
To evaluate this expression we have to compute the next simplified Newton direction $\overline{\delta x_{+}^{\lambda}}$ for the Newton path problem (28) at the new iterate $x_{+}$ for the initial guess $x$. This direction is given by the solution of the equation
${\langle{V_{y(x)}^{-1}}\rangle}^{\prime}(F(x))F^{\prime}(x)\overline{\delta x^{\lambda}_{+}}+V^{-1}_{y(x)}(y(x_{+}))F(x_{+})-(1-\lambda)F(x)=0.$ (33)
These calculations are repeated, in a back-tracking line search, with decreasing damping factor until $\theta_{x_{+}}({x})<\Theta_{acc}$ for some acceptable contraction $\Theta_{acc}$ holds. To achieve this, the damping factor $\lambda$ is readjusted iteratively in the following way. If $\theta_{x_{+}}({x})\geq\Theta_{acc}$, choose $\lambda_{+}\in(0,1]$ such that for user defined parameters $0<\Theta_{des}<\Theta_{acc}$ the inequality $\lambda_{+}\,\theta_{x_{+}}({x})\leq\lambda\,\Theta_{des}$ holds, i.e. we compute the next damping factor by
@@
Once a new iterate $x_{+}=R_{x}(\lambda\delta x)$ is accepted, we do not iterate towards the corresponding point $x(\lambda)$ on the “old” Newton path, but use it as a starting point of a new Newton path, starting at $x_{+}$.
We summarise these ideas in Algorithm 2, which is a modification of [9, Alg. NLEQ-ERR].
@@
and the inner loop terminates after finitely many iterations.
@@
Without going into the details, the proof of the implicit function theorem, which yields local existence of the Newton path (Prop. 6.2), shows that the simplified Newton method converges with a linear rate for sufficiently small choice of $\lambda$. By further reducing $\lambda$, the corresponding rate can be chosen arbitrarily small, in particular faster than $\Theta_{acc}$. Thus, since
[…]
is a lower bound for this rate, we obtain our first statement.
A sufficiently small choice of $\lambda$ is taken within finitely many steps: if $\theta_{R_{x}(\lambda\delta x)}(x)>\Theta_{acc}$, we obtain
@@
Since $\Theta_{des}<\Theta_{acc}$, this yields $\lambda_{+}<C\cdot\lambda\text{ with }C<1$, so in each step, $\lambda$ is reduced about a fixed factor, until the termination criterion is reached. ∎
@@
By the choice of $\Theta_{des}\in(0,1)$ it is possible to adjust how aggressive the step size strategy acts. If $\Theta_{des}\approx 0$, then Newton’s method follows the Newton path very closely and takes short steps. If $\Theta_{des}\approx 1$, then larger steps will be taken, but the methods might perform less robustly for highly nonlinear problems. The choices $\Theta_{des}=0.5$ and $\Theta_{acc}=1.1\Theta_{des}$ work well in practice.